can you help me figure out this linear equation and help me graph it.
y=-2/3x+5

Answer by rfer(16322)

y=-2/3x+5
y=-2/3(0)+5
y=5
(0,5)
--------------
2/3x=-y+5
2/3x=-0+5
2/3x=5x=5*3/2
x=15/2
x=7 1/2
(7 1/2,0)
There is a start for you